OpenStudy (anonymous):

Evaluate the expression: √0.09

myininaya (myininaya):

\[\sqrt{\frac{9}{100}}=\frac{3}{10}\]

OpenStudy (anonymous):

\[.3\] since \[.3^2=.09\]

OpenStudy (anonymous):

actually i like myininaya's reason more than mine

jimthompson5910 (jim_thompson5910):

\[\sqrt{0.09}\] \[\sqrt{\frac{9}{100}}\] \[\frac{\sqrt{9}}{\sqrt{100}}\] \[\frac{3}{10}\] So \[\sqrt{0.09}=\frac{3}{10}=0.3\]

myininaya (myininaya):

:) thanks satellite